### v1

# Perception-as-Control: Fine-grained Controllable Image Animation with 3D-aware Motion Representation

Abstract
Motion-controllable image animation is a fundamental task with a wide range of potential applications. Recent works have made progress in controlling camera or object motion via various motion representations, while they still struggle to support collaborative camera and object motion control with adaptive control granularity. To this end, we introduce 3D-aware motion representation and propose an image animation framework, called Perception-as-Control, to achieve fine-grained collaborative motion control. Specifically, we construct 3D-aware motion representation from a reference image, manipulate it based on interpreted user intentions, and perceive it from different viewpoints. In this way, camera and object motions are transformed into intuitive, consistent visual changes. Then, the proposed framework leverages the perception results as motion control signals, enabling it to support various motion-related video synthesis tasks in a unified and flexible way. Experiments demonstrate the superiority of the proposed framework. For more details and qualitative results, please refer to our project webpage: Perception-as-Control.
Figure 1: Potential applications for Perception-as-Control. By constructing 3D-aware motion representation based on user intentions and utilizing the perception results as motion control signals, the proposed fine-grained motion-controllable image animation framework can be applied to various motion-related video synthesis tasks, such as image-based motion generation (animate image according to user inputs), and video-based motion clone (mimic the entire motions), motion transfer (relocate and scale local motions based on semantic correspondence), motion editing (editing any motion in user-specified regions).
1 Introduction
With the proliferation of generative AI, image animation as a fundamental task is undergoing rapid progress [20, 7, 2, 40, 23, 16, 27, 31]. Various potential application scenarios put forward higher demands for this task, i.e., particularly in achieving precise motion-controllable image animation based on the interpretation of user intentions. Motion-controllable image animation usually takes a single reference image and motion control signals derived from motion representation as input, aiming to generate videos with the expected motions while maintaining the specific appearance. Since rich appearance information is already included in the reference image, the key lies in designing better motion representations and effectively guiding the motion content generation process.
Motion, in general, can be decomposed into camera motion and object motion. Object motion refers to dynamic movements relative to the world coordinate system, while camera motion arises from viewpoint changes. Although researchers have made great progress in designing various motion representations to obtain control signals for camera or object motion control, achieving collaborative camera and object motion control with fine control granularity remains non-trivial. For camera motion control, some methods directly use camera intrinsic and extrinsic parameters [33] or convert them into Plücker coordinates [37, 9, 47] as high-level control signals, while others apply point cloud reconstruction to render incomplete frames [3, 43, 10] or use dense optical flow [23, 15] as more concrete control signals. For object motion control, recent works usually follow the drag-based interaction setting that allows users to draw point trajectories on the reference frame. These methods convert sparse point trajectories into optical flow [33, 16, 45, 15, 23], reorganize point embeddings [27, 34, 5, 4] as control signals. Nevertheless, works [15, 4, 33, 16, 5] focusing on collaborative camera and object motion control still suffer from conflict control issues and insufficient control granularity.
The key observations of this paper are three-fold: 1) Motion representations based on pixel offsets such as optical flow are flexible to modify, while the lack of 3D awareness leads to conflict control issues in local regions with simultaneous camera and object motions. 2) Reconstruction-based motion representations such as point clouds show better spatial consistency, making them effective for camera motion control but less flexible for fine-grained object motion control. 3) For controllable visual content generation tasks that maintain basic spatial structures, diffusion models prefer spatially aligned control signals over high-level ones. Based on these observations, an effective motion representation for fine-grained collaborative motion control should have the following properties. First, the motion representation is equipped with 3D awareness to disentangle local components from the reference image. Second, the motion representation is consistent in space for stable camera motion control and flexible enough for fine-grained object motion control. Third, the motion control signals derived from motion representation are spatially aligned with the generated frames to enhance the connections between them.
Figure 2: 3D-aware motion representation for simplifying and summarizing the 3D scene corresponding to the reference image, which consists of several unit spheres denoting key object parts and a world envelope accommodating them. We move those spheres and observe the scene based on user intentions to transform camera and object motions into intuitive and consistent visual changes.
To this end, we introduce 3D-aware motion representation to facilitate fine-grained collaborative motion control. In controllable image animation applications, users generally have clear intentions about how objects move and how to observe the scene. The intuitive approach is to reconstruct the 3D scene based on the given reference image, drive the objects according to various user intentions, and render the scene from the expected viewpoints. Considering the significant effort required to reconstruct everything in the scene, we notice that human visual attention is predominantly drawn to moving parts, while the remaining parts are generally perceived as a whole [6]. Therefore, as shown in Fig. 2, we simplify and summarize the 3D scene as a combination of several unit spheres representing key object parts (most moving parts and a few static parts) and a vast world envelope akin to The Truman Show to demonstrate how the scene evolves as the observer’s viewpoint changes, i.e., 3D-aware motion representation. We then render the instantiated 3D scene to obtain perception results that reflect camera and object motions and necessary spatial relations.
Based on 3D-aware motion representation, we propose a fine-grained motion-controllable image animation framework, called Perception-as-Control, to achieve collaborative control of arbitrary and fine-grained camera and object motions. The framework takes a single reference image and the perception results of the corresponding 3D-aware motion representation as motion control signals, which are spatially aligned with the reference image. Specifically, we apply two lightweight encoders with similar architectures to encode camera and object control signals separately to avoid RGB-level interference. The encoded control signals are then merged to form the final motion control signals. We employ a diffusion model based on U-Net architecture and inject appearance information from the reference image and motion information from motion control signals, enabling controllable image animation. We introduce a data curation pipeline to prepare training data from in-the-wild videos. Moreover, we propose a three-stage training strategy to balance camera and object motion control and facilitate fine-grained motion control.
Our contributions are three-fold:
  • We introduce 3D-aware motion representation to facilitate fine-grained collaborative motion control, which can be modified and perceived to transform user intentions as spatially-aligned motion control signals.
  • Based on 3D-aware motion representation, we propose a fine-grained motion-controllable image animation framework, which takes advantage of diffusion models for achieving precise and flexible collaborative motion control.
  • The proposed framework can support various motion-related applications via user intention interpretation, such as motion generation, motion clone, motion transfer, and motion editing.
2 Related Work
Camera Motion Control.
Works targeting on camera motion control [33, 37, 47, 9, 3, 43, 7, 29] aim to generate novel view images with spatial consistency. In computer graphics, camera information is typically represented using a set of intrinsic and extrinsic parameters. Methods like MotionCtrl [33] directly repeat these flattened camera parameters to image size as control signals. CameraCtrl [9] and CamCo [37] transform the camera parameters into Plücker coordinates as pixel-wise embedding, resulting in more discriminative camera motion representations. ViewCrafter [43] and I2V-Camera [3] lift the reference image to 3D and render the point cloud to generate incomplete frames. SparseCtrl [7] and DimensionX [29] generate a specific set of training data for each type of camera motion and train individual LoRA models [11] tailored for each camera motion pattern. Concurrent work TrajectoryAttention [36] uses dense pixel trajectories to enhance temporal consistency, which is effective for camera motion control but hard to further support object motion control. Compared to high-level camera motion representations, our 3D-aware motion representation embodies camera motion information via visual changes of the world envelope, which is more intuitive and effective.
Object Motion Control.
Previous works usually follow the drag-based interaction setting for object motion control [42, 16, 27, 23, 45, 34], which conveys user intentions via sparse point trajectories. Methods such as DragNUWA [42], ImageConductor [16], Motion-I2V [27] and MOFA-Video [23] employ optical flow as an intermediate motion representation. The former two methods transform point trajectories into sparse optical flow as control signals, while the latter two methods transform sparse point trajectories into dense optical flows via off-the-shelf or pretrained sparse-to-dense flow generation models. Tora [45] enhances point trajectories into a trajectory map and employs 3D VAE to encode it. DragAnyhting [34] obtains features for each pixel via DDIM inversion [28] and reorganizes those features according to segmentation masks and point trajectories while being less effective in handling motions containing shape variations. Although the above methods support camera control using point trajectories, they suffer from conflict control issues in local regions with simultaneous camera and object motions. Instead of modifying 2D motion representation, we employ a simplified 3D scene and use its inherent perception results as control signals to avoid control conflicts. Due to this design, diffusion models can handle both local object motion and global camera motion simultaneously.
Collaborative Motion Control.
Only a few concurrent works [15, 5, 4] support collaborative motion control. AnimateAnything [15] applies two-stage generation, one for transforming camera and object control into optical flow representation, the other for generating videos based on that. However, they represent camera motion as Plücker coordinates, which leads to imprecise camera motion control. MotionPrompting [5] employs dense point embeddings constructed from 2D point tracking results as motion control signals, while such representation is not friendly enough for complicated motion control. I2VControl [4] also employs dense point embeddings constructed via point transform and projection, while the granularity of object motion control is limited to object segmentation masks. In contrast, we propose Perception-as-Control, which takes motion control signals derived from efficiently constructed 3D-aware motion representation and achieves fine-grained collaborative motion control.
3 Perception-as-Control
Figure 3: Overview. In our data curation pipeline, an in-the-wild video is taken as input, and off-the-shelf visual odometry and 3D point tracking algorithms are applied to obtain camera poses and 3D point locations. Based on these results, 3D-aware motion representation can be constructed and rendered to derive camera and object control signals. During training, those control signals are fed into separate encoders to avoid RGB-level interference and merged afterward. A three-stage training strategy is introduced to facilitate collaborative motion control. During inference, user intentions are first interpreted as 3D-aware motion representation, enabling our framework to support various motion-related applications.
In this section, we first introduce a data curation pipeline for constructing 3D-aware motion representation from in-the-wild videos. Then we describe our fine-grained motion-controllable image animation framework, Perception-as-Control. A three-stage training strategy is introduced to balance camera and object motion control and facilitate fine-grained object motion control.
3.1 Preliminaries
Video Diffusion Model (VDM) adopts the foundational principles of image diffusion probabilistic models, leveraging architectures like denoising U-Net [2, 8] and DiT [40] for video generation. In addition to generating reasonable images, VDM aims to model the temporal dependencies between frames, generating videos that maintain consistent visual content and coherent motion. Let $x_{0}\in\mathbb{R}^{T\times H\times W\times C}$ denote the initial video latent, where $T$ is the number of frames to generate. The forward process of diffusion models is defined as a chain that progressively adds Gaussian noise to the original video, as shown in Eq. 1:
$x_{t}=\sqrt{\hat{\alpha_{t}}x_{t-1}}+\sqrt{{1-\hat{\alpha_{t}}}\epsilon},\epsilon\sim N(0,1),$ (1)
where $t\in{1,\dots,T}$ denotes the diffusion timestep, $\hat{\alpha_{t}}$ regulates the intensity of noise added at each $t$, and $\epsilon$ is drawn from standard Gaussian noise. Thus, the optimization objective of a diffusion model is to predict the probability distribution $p(z_{t-1}|z_{t})$ using a parameterized network $\theta$. The corresponding loss function can be expressed as Eq. 2:
$\mathcal{L}_{\theta}=\mathbb{E}_{x_{0},\epsilon,\mathcal{C},t}[||\epsilon-\hat{\epsilon_{\theta}}(x_{t},\mathcal{C},t)||^{2}_{2}],$ (2)
where $\mathcal{C}$ is an optional parameter denoting the conditions, such as reference images and motion control signals. Typically, VDMs first encode videos into a latent space with the same dimensions as $x_{t}$ to reduce the computational complexity and achieve better results.
3.2 3D-aware Motion Representation
For training purposes, pairs of video data and the corresponding 3D-aware motion representation are required. Unfortunately, commonly used large-scale video datasets predominantly consist of in-the-wild videos lacking camera settings and scene information. One way to recover such information involves estimating camera parameters and reconstructing the scene through monocular 3D reconstruction algorithms such as SfM [46] and then simplifying the reconstructed 3D scene. However, those algorithms are computationally expensive and the scale ambiguity issue leads to non-negligible reconstruction errors. Alternatively, constructing the simplified 3D scene directly via appropriate approximations can bypass these problems and is readily achievable even for in-the-wild videos. Thus, we introduce our data curation pipeline, using off-the-shelf algorithms to construct 3D-aware motion representation for any collectible video, as shown in the top-left part of Fig. 3.
Cut Detection and Motion Filtering.
First, we apply common preprocessing steps to extract video clips containing continuous visual content. We employ a cut detection algorithm and split the original videos into several video clips. Subsequently, we use RAFT [30] to estimate the optical flow between each pair of adjacent frames in a video clip. We then calculate the Frobenius norm of the estimated optical flow as a motion score, which is further used to filter out video clips with non-obvious motion (motion scores lower than $30^{\rm th}$ percentile), the same as previous methods [16]. Next, we construct our 3D-aware motion representation for each video clip. The process is divided into two steps: 3D point tracking to capture local object motion and visual odometry to capture global camera motion.
3D Point Tracking for Local Object Motion.
Key object parts refer to not only moving object parts but also stationary ones, which are critical to revealing spatial relations. They are simplified as unit spheres in the 3D scene, and the number of these spheres is defined according to control granularity flexibly. We require 3D locations of key object parts at each moment to set these spheres in the initial location and simulate their movements. Therefore, we employ SpaTracker [35] to track 2D points in the pixel plane while estimating their depth relative to the camera, i.e., 3D point tracking in each video clip. Note that combining a 2D point tracking algorithm [13] with a depth estimation algorithm [24, 38, 39] has similar functions. In our implementation, we track $25\times 25$ grid points in each video clip, resulting in 3D point tracking results $P=[P^{\rm X},P^{\rm Y},P^{\rm D}]\in\mathbb{R}^{T\times N\times 3}$ ($T$ is the number of frames and $N$ is the number of tracked points).
Visual Odometry for Global Camera Motion.
We employ a world envelope to mark the remaining parts, which is stationary relative to the world coordinate system. To transform global camera motion into visual changes, we can use off-the-shelf algorithms [32, 46] to recover camera trajectories from a sequence of continuous frames and render the world envelope. Despite the scale ambiguity problem, our world envelope, which serves as a scene marker, only requires coarse camera poses instead of extremely precise ones. Thus, we set the camera intrinsic parameters as $I=[W,0,W//2;0,H,H//2;0,0,1]$, where $W$ and $H$ are the width and height of the first frame. In our implementation, we employ TartanVO [32] for this purpose, since it is fast and generalizes well across various cameras. For each video clip, we obtain the estimated camera poses $E=\{E_{t}|t\in\{1,2\dots,T\}\}$, where $T$ is the number of frames, $E_{t}=[R\in\mathbb{R}^{3\times 3},T\in\mathbb{R}^{1\times 3}]$ are camera extrinsic parameters.
Perceptual 3D Scene Rendering.
Once we have obtained the 3D point tracking results $P$ and the estimated camera poses $E$, we construct our 3D-aware motion representation and render it as motion control signals.
First, we set $N$ unit spheres with different colors representing key object parts. These spheres are rendered on the pixel plane according to the camera pose at each timestep and appear as circles that are larger near and smaller far away in the rendering images under perspective projection. Since the estimated depth is not precise enough to determine the absolute 3D positions of these spheres, we normalize point depths $P^{\rm D}$ as $\hat{P}_{\rm d}$ for each video clip to only focus on their relative distance. After that, we set the projected circle radius of the $n^{\rm th}$ sphere to $1-P^{\rm D}_{t,n}$ ($P^{\rm D}_{t,n}$ is the normalized depth of the $n^{\rm th}$ sphere at timestep $t$). For convenience, we generate a continuous color map where each pixel has a distinct color. We then set the color of each sphere to the corresponding color of its projection center in the first frame. Finally, these spheres are projected and rendered as colorful circles with varying radii in the pixel plane.
Then, we construct a cube centered at the origin with the side length of $z_{far}$ ($z_{far}$ is the far clipping plane of a viewing frustum) as an implementation of the world envelope. We apply UV textures in checkerboard style to increase the recognizability of the world envelope, which helps demonstrate how the scene evolves as the viewpoint changes. We align the camera coordinate system to the world coordinate system for the first frame and render the subsequent frames according to the camera pose sequence $E$, without loss of generality. Note that the implementation of the world envelope is not limited to a cube, and its textures can vary beyond the checkerboard style.
To avoid RGB-level interference, we render unit spheres and world envelope separately as two spatial layers, acting as object and camera control signals respectively.
3.3 Network Architecture
We propose a motion-controllable image animation framework called Perception-as-Control based on 3D-aware motion representation, which takes a reference image and a sequence of spatially-aligned motion control signals as input and generates a video as output.
As shown in the right part of Fig. 3, our framework is based on a denoising U-Net architecture, with extra motion modules [8] for modeling temporal information. Under the image animation setting, it is essential to preserve the appearance information contained in the reference image (image control signals) and follow the motion information contained in motion control signals. Previous methods inject spatial-aligned control signals by directly adding them to the input noise [19, 22] or integrating them into denoising U-Net through auxiliary architectures [44, 12, 23], and inject high-level control signals via attention mechanism [41, 21]. Given that the motion control signals are already spatially aligned, in our framework, we inject motion control signals by adding them to the noise and inject the reference image via ReferenceNet [12]. Specifically, we use two lightweight encoders to encode camera control signals and object control signals separately, avoiding RGB-level interference. Then, a fusion module merges the encoded camera control signals and object control signals. We implement the fusion module empirically as a convolution block.
During training, pairs of training data are prepared through our data curation pipeline, and a three-stage training strategy is applied for balance camera and object control signals. During inference, user intentions conveyed in various forms are transformed into 3D-aware motion representation, and the proposed framework is used to generate videos based on the representation.
3.4 Training Strategy
Although the proposed 3D-aware motion representation transforms camera and object motion into a unified form, the training strategy is still important since motions in training data are entangled. To this end, we introduce our three-stage training strategy, which aims to balance camera and object control signals for achieving fine-grained collaborative motion control.
Stage 1: Camera Motion Control Training.
In the first stage, we use video clips from RealEstate10K[48] with camera motion only to train our camera encoder. We temporarily exclude object encoder and fusion module from the framework. Since the rendered spheres can be anywhere in the pixel plane, motion modules in denoising U-Net are trained together. This allows the model to align each rendered sphere with its represented object parts. Thus, we train camera encoder, motion modules in denoising U-Net, and ReferenceNet in this stage.
Stage 2: Collaborative Motion Control Training.
In the second stage, we add video clips from WebVid10M[1] that contain both camera and object motions for training. We add object encoder and fusion module back for collaborative motion control. To speed up the alignment between unit spheres and object parts, we utilize dense unit spheres as object control signals in this stage. We fix camera encoder and train object encoder, fusion module, motion modules, and ReferenceNet in this stage. Moreover, to preserve the effectiveness of both control signals simultaneously, especially in conflict cases, we randomly drop one type of control signals at a fixed dropout rate for video clips from RealEstate10K during training.
Stage 3: Dense-to-Sparse Fine-tuning.
The goal of the third stage is to achieve fine-grained object motion control via sparse unit spheres, which requires the model to determine the control range of each rendered point adaptively.
Considering that object motion tends to occur on salient objects in a video clip, we employ an off-the-shelf segmentation algorithm [17] to obtain salient objects in the reference image and select unit spheres inside the segmentation mask as Set1. Then we calculate the trajectory length of each unit sphere and select those with a trajectory length higher than the $80^{\rm th}$ percentile as Set2. By combining Set1 and Set2, we obtain the dense sphere set and randomly sample $N\in\{1,2,\dots,16\}$ spheres from it as the final sparse sphere set per iteration. In this stage, we fix all components except fusion module and motion modules during training.
In all stages, we optimize our framework with a loss function similar to Eq. 2:
$\mathcal{L}=\mathbb{E}_{x_{0},c_{\rm img},c_{\rm cam},c_{\rm obj},t}[||\epsilon-\epsilon_{\theta}(x_{t},c_{\rm img},c_{\rm cam},c_{\rm obj},t)||^{2}_{2}],$ (3)
where $\epsilon\sim N(0,1)$ is drawn from standard Gaussian noise, $t\in{1,\dots,T}$ denotes the diffusion timestep, and $c_{\rm img}$, $c_{\rm cam}$, $c_{\rm obj}$ are the reference image, camera motion control signals and object motion control signals, respectively.
Figure 4: Case study of camera-only motion control.
4 Experiments
4.1 Experimental Settings
Datasets.
We use video datasets RealEstate10K [48] and WebVid10M [1] for training. RealEstate10K contains videos with only camera motion, and each video is annotated with ground-truth camera intrinsic and extrinsic parameters. WebVid10M is a large-scale dataset containing in-the-wild videos with both camera and object motions, while the quality of these videos is not high enough. After our data curation pipeline, our final training data contains around 6K video clips from RealEstate10K and 35K video clips from WebVid10M.
Implementation Details.
We adopt Stable Diffusion 1.5 (SD1.5) [26] as our base model, and initialize the weights of both denoising U-Net and ReferenceNet using the pretrained weights of SD1.5. We add motion modules the same way as previous works [8, 2] with random initialization. During training, the weights of VAE [14] encoder and decoder, as well as the CLIP [25] image encoder are frozen. During training, we randomly sample 16 frames for each video clip with a resolution of $768\times 512$. We use AdamW [18] as an optimizer. We conduct the training on 8 NVIDIA A100 GPUs. We train around 20k iterations for stage 1, 20k iterations for stage 2, and 50k iterations for stage 3, with a batch size of 1 for convergence.
Figure 5: Case study of object-only motion control.
Figure 6: Case study of collaborative motion control.
4.2 Motion-controllable Image Animation
4.2.1 Camera-only Motion Control
To demonstrate the superiority of Perception-as-Control in camera-only motion control, we select both basic and arbitrary camera motions and visualize them in 3D. As shown in Fig. 4, our framework can generate videos that adhere to the expected camera motion. Owing to the world envelope that shows the direction and magnitude of camera motion and enhances the overall perception, the spatial structures in the generated videos are stable and consistent. Moreover, by adding several unit spheres to essential parts, the generated videos are aware of depth information and can accurately reflect perspective effects under different focal lengths.
Figure 7: Various motion-related applications. Red boxes denote reference images or source videos.
4.2.2 Object-only Motion Control
For object-only motion control, we provide cases in Fig. 5 to demonstrate the superiority of our method in achieving flexible and precise multi-instance and fine-grained motion control. Since there is no camera motion, we project unit sphere centers onto the pixel plane and visualize how they change over time, using colors to represent the direction of movement. As shown in the left part of Fig. 5, we control an increasing number of hot air balloons in the reference image, and the generated videos precisely reflect the movements of each one. In the right part of Fig. 5, when additional controls are applied to the bamboo, the proposed method can adaptively determine the control range for each unit sphere.
4.2.3 Collaborative Motion Control
Our framework also supports collaborative control of both camera and object motions. As illustrated in Fig. 6, we control both camera and object motions and visualize the corresponding 3D-aware motion representation. Our method can handle both large object motions and delicate object motions. The generated videos adhere to the specified camera motion and accurately present adapted object motions.
4.3 Applications
The proposed image animation framework can be used for various motion-related applications by carefully interpreting user intentions, as illustrated in Fig. 1.
Motion Generation refers to image animation based on a reference image and 2D trajectories drawn by users. Motion Clone refers to mimicking both camera and object motions in a source video. Users provide a source video to construct 3D-aware motion representation and edit the first frame as the reference image. Motion Transfer refers to transferring local motions from the source video to the reference image by adaptively relocating and scaling them according to semantic correspondence. Motion Editing has greater control freedom than the above settings. Users can provide segmentation masks and edit motions inside or outside the segmentation masks.
More examples are shown in Fig. 7. For motion generation, we construct 3D-aware motion representation based on the reference image via monouclar depth estimation and manipulate it according to 2D trajectories. The results show that our framework effectively handles various motions including natural swaying, shape changes, intersections, and movements that extend out of the pixel plane. For motion clone, we edit the first frame of each source video as the reference image and construct 3D-aware motion representation from the source video as control signals. The results exactly reproduce the camera and object motions in the source video. For motion transfer, we extract 3D trajectories of unit spheres corresponding to semantic points and relocate them to match the reference image. The results show that local motions from the source video are successfully transferred to objects of different scales at different locations in the reference images. For motion editing, we modify 3D-aware motion representation from the source video based on segmentation masks, e.g., modify several unit sphere trajectories or viewpoints. The results demonstrate the flexibility of our framework.
5 Conclusion
Unlike previous methods that only control camera motion or object motion, we propose a fine-grained controllable image animation framework for collaborative motion control. The framework takes motion control signals derived from 3D-aware motion representation as input, simplifying the original 3D scene to transform camera and object motions into unified and consistent visual changes. Due to the flexibility of the introduced 3D-aware motion representation, our framework can support various motion-related applications by translating user intentions as changes in the representation.